# AniFaceDiff: High-Fidelity Face Reenactment via Facial Parametric Conditioned Diffusion Models

### v1

@@ Abstract
Face reenactment refers to the process of transferring the pose and facial expressions from a reference (driving) video onto a static facial (source) image while maintaining the original identity of the source image. Previous research in this domain has made significant progress by training controllable deep generative models to generate faces based on specific identity, pose and expression conditions. However, the mechanisms used in these methods to control pose and expression often inadvertently introduce identity information from the driving video, while also causing a loss of expression-related details. This paper proposes a new method based on Stable Diffusion [1], called AniFaceDiff, incorporating a new conditioning module for high-fidelity face reenactment. First, we propose an enhanced 2D facial snapshot conditioning approach by facial shape alignment to prevent the inclusion of identity information from the driving video. Then, we introduce an expression adapter conditioning mechanism to address the potential loss of expression-related information. Our approach effectively preserves pose and expression fidelity from the driving video while retaining the identity and fine details of the source image. Through experiments on the VoxCeleb dataset [2], we demonstrate that our method achieves state-of-the-art results in face reenactment, showcasing superior image quality, identity preservation, and expression accuracy, especially for cross-identity scenarios. Considering the ethical concerns surrounding potential misuse, we analyze the implications of our method, evaluate current state-of-the-art deepfake detectors, and identify their shortcomings to guide future research.
@@ 1 Introduction
Face reenactment produces a video of a person engaged in conversation by transferring expressions and poses from a reference (driving) video to a source image. This technique finds applications in various fields, including film production, the video game industry, and video conferencing. Two crucial objectives define this task: 1) preserving the identity and background details of the source image; and 2) transferring pose and expression from the driving video onto the static source image. However, achieving a faithful generation that fulfills these criteria remains a significant challenge.
@@
The advancement of deep generative models, including Generative Adversarial Networks (GANs) [3] and diffusion models [4], has facilitated the successful deployment of convenient and high-quality face reenactment techniques. A typical line of the state-of-the-art GANs or diffusion models based methods [5, 6, 7, 8] relies on flow fields to complete the transferring of the pose and the expression. Initially, keypoints of the source image and the driving video are extracted in an unsupervised manner, followed by the estimation of motion flow fields based on these keypoints to capture pose and expression changes. These flow fields are then utilized to warp either the source image or its features. However, the incorporation of estimated flow fields inadvertently introduces identity information from the driving video onto the source image, resulting in a subpar performance in cases of identity mismatch between the source image and the driving video [9]. Moreover, flow fields often cause distortions, particularly in situations involving significant pose and expression variations between the source image and the driving video.
@@
More recent research on face reenactment has highlighted the potential of pretrained 3D parametric face models [10, 11, 12], for representing the semantic meaning of human faces, including expressions, poses, and identities. Methods based on StyleGAN [13], e.g., StyleHeat [14] and HyperReenact [15], utilize parameters or feature maps from 3D face models to adjust the feature map or weights of GAN-generators. However, these approaches are limited by their reliance on finding a latent representation of the source image within the StyleGAN space that can be manipulated. This often leads to inadequate reconstruction of the source image, including both identity information and background details. Diffusion models have also been extended alongside 3D face models [16, 17] owing to their remarkable performance. However, the conditioning mechanisms of these methods still introduce unexpected identity-related information due to the direct use of the driving image to generate the condition. Furthermore, the typically low resolution of 3D face meshes leads to the loss of necessary expression-related information, particularly mid-frequency details [12].
@@
In this work, we propose a method named AniFaceDiff, for face reenactment by incorporating conditional signals to the pretrained text-to-image Stable Diffusion model [1], as shown in Fig. 1. We introduce a new conditioning mechanism to address the limitations of previous methods. Specifically, 1) to avoid introducing unexpected identity-related information from the driving image, we propose a Facial Shape Alignment (FSA) strategy to form spatial-aligned conditions for the diffusion models. This involves extracting pose and expression parameters from the driving video and shape parameters from the source image using the 3D face model (DECA [12]), and then generating 2D face surface normal snapshots with these facial parameters. 2) To mitigate potential expression-related information loss, we introduce an Expression Adapter (EA) that injects the expression embedding into the cross-attention layers of the denoising UNet [18]. This adapter compensates for the information loss in spatial-aligned conditions, thereby enhancing expression fidelity.
@@
1) We propose a face reenactment framework, AniFaceDiff, based on diffusion models with accurate conditioning. Our approach effectively preserves pose and expression fidelity from the driving video simultaneously retaining the identity and background details of the source image in a one-shot setting (solely utilizing a single source image).
2) We design a new conditioning mechanism based on 3D face model, which furnishes both accurate spatial and non-spatial conditions for diffusion models, thereby facilitating faithful reenactment.
3) Experiments on the VoxCeleb [2] benchmark demonstrate that our method achieves state-of-the-art performance in face reenactment tasks. Furthermore, our method has robust generalization capabilities, even when applied to out-of-domain data.
Figure 1: The overview of the proposed method. A VAE [19] encoder and the ReferenceNet extract detailed features from the source image, which are then merged into the Stable Diffusion backbone Denoising UNet via spatial-attention. A CLIP Image Encoder [20] extracts semantic features from the source image, which is then injected into both the ReferenceNet and the Denoising UNet via cross-attention. The pose and expression conditioning is portrayed within the red dashed box. The facial shape of the source image and the pose and expression of the driving video are extracted to form 2D surface normal facial snapshots through Facial Shape Alignment. These snapshots are then encoded and concatenated with the input of the Denoising UNet. The expression of driving frames is further injected into the cross-attention layers of the Denoising UNet by the expression adapter to improve the expression consistency. Temporal-attention aims to improve the temporal consistency across frames. After iterative denoising, the output of the Denoising UNet is decoded into the final animated video by a VAE decoder.
@@ 2 Related Work
Neural face reenactment
The development of deep generative models has facilitated convenient and high-quality face reenactment applications. A prevalent approach involves warping the source image or latent representation using deformation fields (e.g., optical flow fields) to convey pose and transfer expressions [21, 22, 5, 6, 9, 7, 23, 24, 25, 26, 27, 8]. For instance, Monkey-Net [22] employs a motion transfer network alongside unsupervised keypoint detection and dense motion prediction to generate animated frames. FOMM [5] enhances Monkey-Net by incorporating motion field computation with a first-order Taylor expansion approximation, encompassing keypoints and affine transformations. Face vid2vid [6] extends FOMM for free-view talking face video generation utilizing a free-view keypoints representation. TPSMM [7] introduces thin-plate spline motion estimation to generate a more flexible optical flow. However, optical flow fields are susceptible to introducing artifacts and blur, especially in instances of large identity or pose mismatches between the source image and driving video. To mitigate this issue, MRAA [9] proposes an Animation Via Disentanglement (AVD) network to separate the control of shape and pose. However, improvements in results are not as significant for objects like faces.
Another line of research exploits the learned prior of pretrained 3D parametric face models. PIRenderer [28] utilizes the parameters of 3D morphable face models (3DMMs) to adjust intermediate results from the warping network. HeadGAN [29] generates faces with warped source features, 3d face shape and audio. StyleHeat [14], based on StyleGAN [13], integrates 3DMMs with StyleGAN by predicting flow fields with 3DMM parameters and warping the feature map from the encoder of GAN inversion. However, these methods based on 3D face models still depend on flow fields and often necessitate additional refinement to attain satisfactory results. In addition, StyleGAN-based methods rely on GAN inversion, which may struggle to achieve high-fidelity reconstruction following extensive edits [15]. HyperReenact [15] employs the feature map of the pretrained DECA encoder to update the weights of the pretrained StyleGAN generator, allowing it to perform well even under large pose variations. PASL [30] proposes pose adapted shape learning for large pose reenactment. Our method aims to enhance the accuracy of both identity preservation and expression.
@@
Conditioning serves as a crucial element in maximizing the potent generative capabilities for various downstream tasks, enabling controllable generation [39, 40, 16, 17, 41, 42, 43]. The two primary conditioning mechanisms for diffusion models involve concatenation and cross-attention [1]. Liu et al. [44] propose an alternative intuitive approach based on [32] and a CLIP-based encoder for semantic diffusion guidance in both text and image conditioning. ControlNet [45] is extensively utilized for spatial-aligned conditioning by updating each layer of the Stable Diffusion backbone UNet via trainable copy and zero convolution layers. IP-adapter [46] integrates image conditions into pretrained text-to-image diffusion models using a decoupled cross-attention mechanism. Our pose and expression conditioning module employs designs similar to those of ControlNet and IP-adapter.
Human image animation with diffusion models
Animating a static image into a temporally consistent video [47, 48, 49, 50, 51] has garnered significant attention among researchers. In human image animation, DisCo [52] introduces human appearance conditioning to Stable Diffusion using a CLIP image encoder and separate ControlNets for background and pose conditions to generate human dancing videos. Animate Anyone [53] employs spatial attention and a ReferenceNet, a copy of UNet, to incorporate detailed information into Stable Diffusion and demonstrate its generalization capability. Champ [54] enhances previous approaches by utilizing a 3D human parametric model with multiple pose conditions. In audio-driven talking head generation, EMO [55] proposes a direct audio-to-video framework leveraging pretrained wave2vec [56]. VLOGGER [57] predicts face and body parameters from audio to render dense masks as conditions for generating talking avatars. We are inspired by previous efforts utilizing the ReferenceNet and spatial-aligned pose conditions due to the similarity of the task. We focus on achieving accurate face reenactment guided by pose and expression while minimizing the impact of identity mismatches.
@@ 3 Method
Our method is designed for one-shot face reenactment, which involves generating a realistic video based on the guidance of a single source image and a driving video. The framework of the proposed method is illustrated in Fig. 1. First, we give a brief introduction of the foundational model – Stable Diffusion and the 3D face model (Sec.3.1). Second, we introduce the overall structure of the framework (Sec.3.2). Then, we detail the proposed pose and expression conditioning mechanism (Sec.3.3). Finally, we illustrate the detailed information of training and inference in Sec.3.4
@@ 3.1 Preliminaries
3D morphable face models (3DMMs). 3DMMs are parametric models designed to accurately represent facial shape and expression. These models are constructed using dimensionality reduction techniques such as principal component analysis (PCA) and are capable of reconstructing 3D faces based on 2D images. FLAME [11] is a notable example of a 3DMM that relies on standard vertex-based linear blend skinning with blendshapes to generate a mesh with 5023 vertices, which is formulated as:
@@
where $\beta$, $\theta$, and $\psi$ represent parameters of identity, pose, and expression, respectively. $W$ is the blend skinning function that rotates the vertices in $T_{p}$ around joints J smoothed by blendweights $\mathcal{W}$. Detailed information can be found in FLAME [11]. In this paper, we utilize the widely used 3D face model DECA [12] which is capable of estimating parameters, including those of the FLAME model, from a single image. This allows for the reconstruction of a 3D face with detailed facial geometry.
@@ 3.2 Framework Architecture
In this section, we illustrate the overall framework of our method, which takes a source image and a driving video as inputs and outputs a reenacted video. We formulate face reenactment as conditional generation, where all the identity, pose, and expression of the generated content are controllable via the corresponding conditioning mechanism. Our method is based on the pretrained Stable Diffusion 1.5 and follows a similar structure to the backbone denoising UNet. First, the source image is encoded into a latent space using a VAE encoder, and its features are extracted by a ReferenceNet. Second, we utilize a CLIP image encoder to extract semantic information, which is crucial for preserving the identity of the source image. This semantic information is injected into the ReferenceNet and Denoising UNet via cross-attention. To effectively introduce pose and expression, we propose a pose and expression conditioning module as elaborated in 3.3. Furthermore, we employ Temporal Modules to generate temporally consistent content as we detail below. The diffusion model performs iterative denoising on the latent space and finally transforms the denoised output back to the pixel space through a VAE decoder to get the generated video.
ReferenceNet. ReferenceNet is a copy of the backbone denoising UNet (without temporal modules) used to extract features at multiple resolutions containing the face and background of the source image. It has been widely utilized [53, 55, 54] to improve the appearance consistency between the reference and the output. These features are then merged into the denoising UNet using a spatial-attention mechanism similar to Animate Anyone [53]. Specifically, the output of each self-attention layer of ReferenceNet ($z_{1}\in\mathbb{R}^{b\times c\times h\times w}$) is repeated $\mathit{f}$ times (the length of the video clip) and concatenated with that of the denoising UNet ($z_{2}\in\mathbb{R}^{b\times c\times f\times h\times w}$). The model then applies self-attention on the concatenated output ($z_{concat}\in\mathbb{R}^{b\times c\times f\times h\times 2w}$) and takes the first half of the feature map as the final output ($z_{out}\in\mathbb{R}^{b\times c\times f\times h\times w}$).

### v2

# AniFaceDiff: Animating Stylized Avatars via Parametric Conditioned Diffusion Models

@@ Abstract
Animating stylized avatars with dynamic poses and expressions has attracted increasing attention for its broad range of applications. Previous research has made significant progress by training controllable generative models to synthesize animations based on reference characteristics, pose, and expression conditions. However, the mechanisms used in these methods to control pose and expression often inadvertently introduce unintended features from the target motion, while also causing a loss of expression-related details, particularly when applied to stylized animation. This paper proposes a new method based on Stable Diffusion [1], called AniFaceDiff, incorporating a new conditioning module for animating stylized avatars. First, we propose a refined spatial conditioning approach by Facial Alignment to prevent the inclusion of identity characteristics from the target motion. Then, we introduce an Expression Adapter that incorporates additional cross-attention layers to address the potential loss of expression-related information. Our approach effectively preserves pose and expression from the target video while maintaining input image consistency. Extensive experiments demonstrate that our method achieves state-of-the-art results, showcasing superior image quality, preservation of reference features, and expression accuracy, particularly for out-of-domain animation across diverse styles, highlighting its versatility and strong generalization capabilities. This work aims to enhance the quality of virtual stylized animation for positive applications. To promote responsible use in virtual environments, we contribute to the advancement of detection for generative content by evaluating state-of-the-art detectors, highlighting potential areas for improvement, and suggesting solutions.
Figure 1: Stylized animation results produced by our method. In the visualizations, the first columns are reference images and the following three columns are animated results. Our method generates high-quality results using various reference and target motion images that differ in pose, expression, and even domain. In particular, it effectively preserves the reference appearance and style with the pose and expression of the target motion, even under significant pose variations. Furthermore, it demonstrates a strong capability to handle avatars across a wide range of styles and domains. We aim on virtual stylized character animation. All characters in this paper are generated by [2] and do not correspond to real individuals. More examples can be found in Fig. 6.
@@ 1 Introduction
The animation of stylized avatars refers to the process of animating static, stylized characters by poses and expressions derived from target motion. This technique has a wide range of applications for film and animation production, video conferencing, customer service, and privacy protection. For instance, it can be utilized to safeguard patients’ appearance privacy in telemedicine and to enhance engagement in virtual learning through artistic representations. Two crucial objectives define this task: 1) preserving the appearance and background details of the stylized reference image; and 2) accurately reflecting the pose and expressions from the target motion. However, achieving high-quality stylized animation that meets these objectives remains a significant challenge.
@@
The advancement of deep generative models, including Generative Adversarial Networks (GANs) [3] and diffusion models [4], has facilitated the success of high-quality avatar animation. A typical line of the state-of-the-art GANs or diffusion models based methods [5, 6, 7, 8] relies on flow fields to complete the transferring of the pose and the expression. Initially, keypoints of the inputs are extracted in an unsupervised manner, followed by the estimation of motion flow fields based on these keypoints to capture pose and expression changes. These flow fields are then utilized to warp either the reference image or its features. However, the incorporation of estimated flow fields inadvertently introduces identity characteristics from the motion video onto the reference image, resulting in a subpar performance in cases of identity mismatch between the reference image and the target video [9]. This issue is particularly pronounced in stylized animation, where the appearance inconsistency can be more significant, further exacerbating the problem. Moreover, flow fields often cause distortions, particularly in situations involving significant pose and expression variations between the reference image and the target motion.
@@
More recent research has highlighted the potential of pretrained 3D Morphable Face Models (3DMMs) [10, 11, 12], for representing the semantic meaning of faces, including expressions, poses, and other features. Methods based on StyleGAN [13], such as StyleHeat [14], utilize parameters or feature maps from 3D morphable models to adjust the feature map or weights of GAN-generators. However, these approaches are limited by their reliance on finding a latent representation of the reference image within the StyleGAN space that can be manipulated. This often leads to inadequate reconstruction of the reference image, including both appearance and background details. Diffusion models have also been extended alongside 3D morphable models [15, 16] owing to their remarkable performance. However, the conditioning mechanisms of these methods still introduce unexpected identity-related information due to the direct use of the target motion to generate the condition. Furthermore, the typically low resolution of 3D face meshes leads to the loss of necessary expression-related information, particularly mid-frequency details [12], which are critical for achieving expressive animation in stylized avatars.
@@
In this work, we propose a method named AniFaceDiff, for stylized and diverse avatar animation by incorporating conditional signals to the pretrained text-to-image Stable Diffusion model [1]. We introduce a new conditioning mechanism to address the limitations of previous methods. Specifically, 1) to avoid introducing unexpected identity-related information from the target motion, we propose a Facial Alignment (FA) strategy to form spatial-aligned conditions for the diffusion models. This process involves extracting pose and expression parameters from the target video, as well as shape parameters from the reference image, using the 3D morphable model (DECA [12]). These parameters are used to generate 2D surface normal snapshots, which are then encoded as part of the input for the denoising UNet. [17]. 2) To compensate for the expression-related information loss in spatial-aligned conditions, we introduce an Expression Adapter (EA) that integrates the expression embedding from the target video into the denoising UNet using additional cross-attention layers, which are combined with the cross-attention layers for the CLIP embedding [18] of the reference image. As shown in Fig. 1, our results demonstrate the effectiveness of our methods in handling avatars with diverse styles and generating high-fidelity outputs. Furthermore, we aim to improve virtual stylized animation for positive applications. To promote responsible development, we contribute to the improvement of detection for generative content, reinforcing safeguards against potential misuse.
@@
1) We propose a framework based on diffusion models, incorporating an accurate conditioning approach for animation across diverse styles. Our approach effectively preserves pose and expression fidelity from the target motion while maintaining the appearance and background details of the stylized reference image.
2) We design a new conditioning mechanism based on 3D morphable models, which furnishes both accurate spatial and non-spatial conditions for diffusion models, thereby facilitating faithful animation.
3) Extensive experiments demonstrate that our method achieves state-of-the-art performance in both out-of-domain and within-domain animation.
4) We evaluate the generalization capabilities of state-of-the-art synthetic face detectors, helping to identify potential weaknesses and propose possible solutions.
Figure 2: The overview of the proposed method. A reference image and a target video serve as inputs. The expression from the target frames and the CLIP reference embedding are integrated into the decoupled cross-attention layers of the Denoising UNet via the expression adapter, enhancing expression consistency while preserving the semantic information from the reference image. The head shape from the reference image, along with the pose and expression from the target motion, are extracted and scaled to generate 2D surface normal snapshots through Facial Alignment. These snapshots are subsequently encoded and incorporated into the input of the Denoising UNet. The ReferenceNet plays a key role in preserving detailed features from the reference image, while temporal attention focuses on maintaining temporal consistency across frames. After iterative denoising, the final output from the Denoising UNet is decoded into the resulting animated video.
@@ 2 Related Work
Motion-guided avatar animation
The development of deep generative models has facilitated high-quality creation of avatars. A prevalent approach involves warping the reference image or latent representation using deformation fields (e.g., optical flow fields) to convey pose and transfer expressions [19, 5, 6, 9, 7, 20, 21, 22, 23, 8]. For instance, FOMM [5] enhances Monkey-Net by incorporating motion field computation with a first-order Taylor expansion approximation, encompassing keypoints and affine transformations. Face vid2vid [6] extends FOMM for free-view talking head video generation utilizing a free-view keypoints representation. TPSMM [7] introduces thin-plate spline motion estimation to generate a more flexible optical flow. However, optical flow fields are susceptible to introducing artifacts and blur, especially in instances of large pose mismatches between the reference image and target video. To mitigate this issue, MRAA [9] proposes an Animation Via Disentanglement (AVD) network to separate the control of shape and pose. However, improvements in results are not as significant for objects like faces.
Another line of research exploits the learned prior of pretrained 3D parametric morphable models. PIRenderer [24] utilizes the parameters of 3D morphable models to adjust intermediate results from the warping network. HeadGAN [25] generates heads with warped reference features, 3d face shape and audio. StyleHeat [14], based on StyleGAN [13], integrates 3DMMs with StyleGAN by predicting flow fields with 3DMM parameters and warping the feature map from the encoder of GAN inversion. However, these methods based on 3D morphable models still depend on flow fields and often necessitate additional refinement to attain satisfactory results. In addition, StyleGAN-based methods rely on GAN inversion, which may struggle to achieve high-fidelity reconstruction following extensive edits [26]. HyperReenact [26] employs the feature map of the pretrained DECA encoder to update the weights of the pretrained StyleGAN generator, allowing it to perform well even under large pose variations. PASL [27], on the other hand, employs pose-adapted shape learning specifically designed for handling significant pose variations. Our method builds on diffusion models and aims to enhance both the preservation of stylized appearance and the accuracy of expression through a carefully designed conditioning module, particularly for avatars with diverse styles.
@@
Conditioning serves as a crucial element in maximizing the potent generative capabilities for various downstream tasks, enabling controllable generation [36, 37, 15, 16, 38, 39, 40]. The two primary conditioning mechanisms for diffusion models involve concatenation and cross-attention [1]. Liu et al. [41] propose an alternative intuitive approach based on [29] and a CLIP-based encoder for semantic diffusion guidance in both text and image conditioning. ControlNet [42] is extensively utilized for spatial-aligned conditioning by updating each layer of the Stable Diffusion backbone UNet via trainable copy and zero convolution layers. IP-Adapter [43] integrates image conditions into pretrained text-to-image diffusion models using a decoupled cross-attention mechanism. Our pose and expression conditioning module is inspired by ControlNet and IP-Adapter.
Character animation with diffusion models
Animating a static image into a temporally consistent video [44, 45, 46, 47, 48] has garnered significant attention among researchers. DisCo [49] introduces appearance conditioning to Stable Diffusion using a CLIP image encoder and separate ControlNets for background and pose conditions to generate dancing videos. Animate Anyone [50] employs spatial attention and a ReferenceNet, a copy of UNet, to incorporate detailed information into Stable Diffusion and demonstrate its generalization capability. Champ [51] enhances previous approaches by utilizing a 3D human parametric model with multiple pose conditions. In audio-driven talking head generation, EMO [52] proposes a direct audio-to-video framework leveraging pretrained wave2vec [53]. VLOGGER [54] predicts face and body parameters from audio to render dense masks as conditions for generating talking avatars. We are inspired by previous efforts utilizing the ReferenceNet and spatial-aligned pose conditions due to the similarity of the task. Our approach focuses on stylized avatar animation using diffusion models, minimizing identity mismatches through Facial Alignment with shape and scale alignment, and mitigating expression loss using the Expression Adapter.
@@ 3 Method
Our method is designed for one-shot animation across diverse styles, which involves generating a video based on the guidance of a single reference image and target motion. The framework of the proposed method is illustrated in Fig. 2. First, we give a brief introduction of the foundational model – Stable Diffusion and the 3D morphable models (Sec.3.1). Second, we introduce the overall structure of the framework (Sec.3.2). Then, we detail the proposed pose and expression conditioning mechanism (Sec.3.3). Finally, we illustrate the detailed information of training and inference in Sec.3.4
@@ 3.1 Preliminaries
3D morphable face models (3DMMs). 3DMMs are parametric models designed to accurately represent head shape and expression. These models are constructed using dimensionality reduction techniques such as principal component analysis (PCA) and are capable of reconstructing 3D faces based on 2D images. FLAME [11] is a notable example of a 3DMM that relies on standard vertex-based linear blend skinning with blendshapes to generate a mesh with 5023 vertices, which is formulated as:
@@
where $\beta$, $\theta$, and $\psi$ represent parameters of shape, pose, and expression, respectively. $W$ is the blend skinning function that rotates the vertices in $T_{p}$ around joints J smoothed by blendweights $\mathcal{W}$. Detailed information can be found in FLAME [11]. In this paper, we utilize the widely used 3D morphable model DECA [12] which is capable of estimating parameters, including those of the FLAME model, from a single image. This allows for the reconstruction of a 3D head with detailed head geometry.
@@ 3.2 Framework Architecture
In this section, we illustrate the overall framework of our method, which takes a reference image and a target video as inputs and outputs an animated video. We formulate avatar animation as conditional generation, where all the appearance, pose, and expression of the generated content are controllable via the corresponding conditioning mechanism. Our method is based on the pretrained Stable Diffusion 1.5 and follows a similar structure to the backbone denoising UNet. First, the reference image is encoded into a latent space using a VAE encoder [56], and its features are extracted by a ReferenceNet. Second, we utilize a CLIP image encoder to extract semantic information, which is crucial for preserving the characteristics of the reference image. This semantic information is injected into the ReferenceNet and Denoising UNet via cross-attention. To effectively introduce pose and expression, we propose a pose and expression conditioning module as elaborated in 3.3. Furthermore, we employ Temporal Modules to generate temporally consistent content as we detail below. The diffusion model performs iterative denoising on the latent space and finally transforms the denoised output back to the pixel space through a VAE decoder to get the generated video.
ReferenceNet. ReferenceNet is a copy of the backbone denoising UNet (without temporal modules) used to extract features at multiple resolutions containing the head, style, and background of the reference image. It has been widely utilized [50, 52, 51] to improve the appearance consistency between the reference and the output. These features are then merged into the denoising UNet using a spatial-attention mechanism similar to Animate Anyone [50]. Specifically, the output of each self-attention layer of ReferenceNet ($z_{1}\in\mathbb{R}^{b\times c\times h\times w}$) is repeated $\mathit{f}$ times (the length of the video clip) and concatenated with that of the denoising UNet ($z_{2}\in\mathbb{R}^{b\times c\times f\times h\times w}$). The model then applies self-attention on the concatenated output ($z_{concat}\in\mathbb{R}^{b\times c\times f\times h\times 2w}$) and takes the first half of the feature map as the final output ($z_{out}\in\mathbb{R}^{b\times c\times f\times h\times w}$).